Find dy/dx of the curve y^2=7+2xy
Answer by stanbon(75887)

Find dy/dx of the curve
y^2=7+2xy
-------
2yy' = 0 + 2[xy' + y]
----
2yy' = 2xy' + 2y
----
yy' = xy' + 2y
(y-x)y' = 2y
---
y' = (2y)/(y-x)
======================
Cheers,
Stan H.